(7/(5x-10))-(3/20)=(7/(4x-8))
please solve I got outrageous numbers
my answer was:
x=(30/36)+or-the square root of (707/36)i

Answer by stanbon(48545)

(7/(5x-10))-(3/20)=(7/(4x-8))
Factor the denominators:
-------------------
7/[5(x-2)]] -[3/20] = 7/[4(x-2)]
----------
Multiply thru by 20(x-2):
7(4) - 3(x-2) = 7(5)
-----------------
28 -3x+6 = 35
3x = -1
x = -1/3
==================
Cheers,
Stan H.